# RDF Stream Taxonomy: Systematizing RDF Stream Types in Research and Practice

### v2

@@ 2.6 Summary
The review uncovered a wide variety of task formulations for RDF streams. It should be noted that most of the identified works use the same term (RDF stream) to describe very different concepts. The review also highlighted an important distinction between theoretical models of RDF streams (e.g., timestamped triples) and practical ones. This distinction was also observed earlier in several works [5, 8, 12, 49]. Another contrasting feature is the use of graphs or datasets as stream elements, versus triple or quad statements. This was also noticed in earlier works focusing on the practical aspects of implementing RDF streams [5, 29].
Worth noting is that some works have several views of the same RDF stream. For example, ERI [17] views its RDF streams as a flow of triples, blocks (RDF graphs), or subject-molecules (RDF subgraphs), depending on the level of abstraction. This suggests that describing the type of an RDF stream depends on the context in which it is considered, or, metaphorically, one’s perspective. In fact, an N-Triples file would be considered by RDF4J’s Rio [16] as a stream of triples, while an application like DBpedia-Live [44] or IPSM [20] would see it as a single element in a stream of graphs.
@@ 3 Proposed Systematization
In what follows, we propose a set of definitions for the types of RDF streams present in the literature and current practice. Then, Section 3.2 organizes these definitions in the RDF Stream Taxonomy, and Section 3.3 links the proposed stream types to the state of the art. The basis for all following definitions is the RDF 1.1 W3C Recommendation [10], with the aim of making the definitions readily applicable in the context of existing, RDF 1.1-compliant software.
The model-theoric semantics of RDF streams are not considered in this work. Instead, the proposed semi-formal definitions can be used as a universal basis on top of which the formal semantics can be defined, as needed. This is motivated by, firstly, the semantics of RDF streams being irrelevant for some of the discussed application areas (e.g., streaming protocols, streaming I/O). Secondly, only the semantics of RDF graphs are well-defined – there are no agreed-upon formal semantics for RDF datasets [66], and there are similarly many ways to approach this topic for RDF streams. Therefore, we leave this for future discussion.
@@ 3.1 RDF Stream Definitions
Definition 1
An RDF stream is an ordered, potentially infinite sequence of RDF stream elements.
This definition by itself cannot be used in practice, as it does not specify what are the RDF stream elements. Thus, we specialize this definition further:
Definition 2
@@
Definition 3
A flat RDF stream is an RDF stream whose elements are statements (either RDF triples or RDF quads).
Grouped and flat RDF streams are still abstract, however, they correspond to an important distinction between streams of groups of statements and streams of statements. With this, we can define the concrete types of RDF streams:
Definition 4
An RDF graph stream is a grouped RDF stream whose elements are unnamed (default) RDF graphs.
In other words, an element of an RDF graph stream is a set of triples. This is a commonly used formulation that can be further specialized:
Definition 5
An RDF subject graph stream is an RDF graph stream in which every element contains an IRI node (called the subject node) that uniquely identifies the graph in the stream. Every other node in the graph can be reached by traversing triples, starting from the subject node.
RDF subject graph streams or similar formulations were proposed to handle timestamped RDF streams (TA-RDF [48] and Stream Containers [49]). ERI [17] uses a similar formulation for subject-molecules, but its subject IRIs are unique only within one block. The proposed definition is slightly broader than the timestamped models, while still being useful – the advantage is that the element can be identified by an IRI, without using RDF datasets. The rest of the element can be discovered with what can be intuitively thought of as RDF graph connectivity.
@@
Definition 6
An RDF dataset stream is a grouped RDF stream whose elements are RDF datasets.
It should be noted here that although an RDF graph stream may seem like a special case of an RDF dataset stream, it is not. According to RDF 1.1, an RDF dataset is not a generalization of an RDF graph, but rather a collection of RDF graphs. Thus, the definitions of RDF dataset streams and RDF graph streams are not directly related in the taxonomy. It should be noted, however, that an RDF graph stream can be trivially transformed into an RDF dataset stream.
@@
Definition 7
An RDF named graph stream is an RDF dataset stream in which every element has exactly one named RDF graph pair $\langle n,G\rangle$, where $G$ is an RDF graph, and $n$ is the graph name. Apart from graph $G$, the dataset may contain any number of triples in the default graph.
This is a simplified version of the RSP Data Model [47], without the temporal aspect. Although we could find only one reference to this stream type in the literature (RSP engines can consume such streams [47]), it is an intuitively important specialization of the RDF dataset stream. Here, every element corresponds to one named RDF graph with some optional information (metadata) about it. As with RDF subject graph streams, this has the advantage of identifying an element by IRI, albeit with more flexibility as to the element’s contents.
@@
Definition 9
A timestamped RDF named graph stream is an RDF named graph stream in which every element is a timestamped named graph. The elements that share the same timestamp predicate $p$ are ordered by the partial order associated with $p$.
These definitions were derived directly from the draft RSP Data Model [47].
Finally, we define concrete flat RDF stream types:
Definition 10
@@
Definition 11
A flat RDF quad stream is a flat RDF stream whose elements are quads.
An “RDF quad” is a term that is often used by practitioners, but not defined in the RDF 1.1 Recommendation. It is however defined in the RDF 1.2 Working Draft [25], and thus we use the RDF 1.2 definition of quads here.
@@
Grouped and flat RDF streams are notably interrelated, as any RDF graph stream can be flattened into an RDF triple stream, by simply listing and concatenating the statements in each of its elements – the same is true for RDF dataset streams and RDF quad streams. The reverse is also possible, the elements of a flat stream can be grouped into a grouped stream. Such operations are in fact common in streaming applications [5, 17, 29, 50].
Finally, for all of the above definitions, their generalized variants can be trivially derived (e.g., generalized RDF graph stream), by using generalized triples and generalized datasets. We skip these definitions for the sake of brevity.
@@ 3.2 RDF Stream Taxonomy
The proposed definitions form a taxonomical structure of task formulations (RDF Stream Taxonomy – RDF-STaX), summarized in Figure 1. This overview allows us to observe that each concrete RDF stream type uses one of the four element types: graph, dataset, triple, and quad, possibly with additional restrictions on them. In fact, these four types of elements seem to be the only conceivable ones that can be derived from the RDF 1.1 specification. For each of them, representing the element is a known and solved problem (RDF serializations), and thus they can be easily used in practical applications.
@@
3.3 Correspondence to the State of the Art
Finally, we link the proposed RDF stream types to the definitions used in state-of-the-art research and software – Table 1 presents such a summary. As noted in Section 2.6, a single RDF stream can be viewed from multiple perspectives, which is reflected in the table, with some items appearing more than once. In some cases the correspondence is not ideal and there may be insignificant differences between the definitions proposed here and those in the literature. The short descriptions presented in the table are expanded upon in the nanopublications dataset, as described in Section 5.1.
@@ 3.3 Correspondence to the State of the Art
RDF graph stream
DBpedia-Live [44], ERI (blocks) [17], GoT [36], IPSM (logical level) [20], Jelly (stream frames) [50], Keskisärkkä & Blomqvist [29], Live open scientific knowledge graphs [37], RDF EXI [28], RDSZ [18], RSP Data Model (can be consumed) [47], RSP4J YASPER (can be consumed) [57], S-HDT [26], VoCaLS [59]
@@
ERI (subject-molecule stream, only within a block) [17], TA-RDF [48], Stream Containers [49]
RDF dataset stream
IPSM (output) [20], Jelly (stream frames) [50], RMLStreamer [24, 45], RSP Data Model (can be consumed) [47], Semantic Annotation enabler [55]
@@
RSP Data Model (can be consumed) [47]
@@
Linked Data Notifications for RDF streams [7], RSP Data Model [47], Tappolet & Bernstein [56], TripleWave [41], Wu et al. [63]
@@
Apache Jena RIOT [2], ERI (high level) [17], Groppe et al. [23], Jelly (high level) [50], Keskisärkkä & Blomqvist (mentioned) [29], RDF4J Rio [16], RSP4J YASPER (can be consumed) [57], SPARQL-Generate [38], VoCaLS [59]
@@
Apache Jena RIOT [2], CARML [9], Jelly (high level) [50], RDF4J Rio [16]
@@ 4 RDF-STaX Ontology
To make the taxonomy readily applicable, it was modeled as an OWL 2 DL ontology [^3] https://w3id.org/stax/1.0.1/ontology (Figure 2). Stream types are instances of either stax:AbstractStreamType or stax:ConcreteStreamType class, with taxonomical relations being realized using SKOS in the OWL DL version [27, 42]. Each stream type is accompanied by a label, formal definition, and usage example.
@@
The ontology includes rich semantic relations between stream types. Properties stax:canBeFlattenedInto and stax:canBeGroupedInto relate grouped and flat stream types that can be easily converted into one another (e.g., an RDF graph stream can be flattened into a flat RDF triple stream). The property stax:canBeTriviallyExtendedInto is used for graph- and triple-based streams, to indicate that they can be trivially turned into dataset- and quad-based streams, respectively. This transformation consists of stating explicitly that the triples are in the default graph. Object property chains are defined in the ontology to allow reasoning on these relations, exploiting the taxonomical relations from SKOS.
@@
It should be stressed that RDF-STaX does not define a class for RDF streams, but rather classes for RDF stream types. It also defines a class for RDF stream type usages, which can be thought of as subjective views of how an RDF stream is used. As described in detail in the documentation [^4] https://w3id.org/stax/use-it, the expected usage pattern is for users to create instances of the stax:RdfStreamTypeUsage class to annotate research works, software, or datasets, with each instance corresponding to a different view or use case of the stream. This pattern was selected due to the inherent subjectivity of determining the types of RDF streams, as discussed in Section 2.6. Below is an example of how this pattern may be used to annotate a DCAT [6] Dataset:
@prefix dcat: <http://www.w3.org/ns/dcat#> .

### v3

@@ 2.6 Summary
The review uncovered a wide variety of task formulations for RDF streams. It should be noted that most of the identified works use the exact same term (RDF stream) to describe very different concepts. The review also highlighted an important distinction between theoretical models of RDF streams (e.g., timestamped triples) and practical ones, making it clear that direct comparisons between the two are not possible. This distinction was also observed earlier in several works Bonte and Tommasini (2023); Calbimonte and Corcho (2016); Dell’Aglio et al. (2013); Schraudner and Harth (2022). Another contrasting feature is the use of graphs or datasets as stream elements, versus singular triple or quad statements. This was also noticed in earlier works focusing on the practical aspects of implementing RDF streams Bonte and Tommasini (2023); Keskisärkkä and Blomqvist (2013).
Worth noting is that some works have several views of the same RDF stream. For example, ERI Fernández et al. (2014) views its RDF streams as a flow of triples, blocks (RDF graphs), or subject-molecules (RDF subgraphs), depending on the level of abstraction. This suggests that describing the type of an RDF stream depends on the context in which it is considered, or, metaphorically, one’s perspective. In fact, an N-Triples file would be considered by RDF4J’s Rio Eclipse Foundation, Inc. (2023) as a stream of triples, while an application like DBpedia-Live Morsey et al. (2012) or IPSM Ganzha et al. (2017) would see it as a single element in a stream of graphs.
The diversity of problem statements encountered in the literature clearly suggests that expecting there to be a single model of RDF streams may be unrealistic, especially given that each stream may be understood in several different ways, depending on the perspective or context. Each of the identified approaches is motivated by some use case and thus is valid. Therefore, we must seek solutions to better understand and harmonize this task diversity, which is the primary goal of RDF-STaX.
@@ 3 Proposed Systematization
In what follows, we propose a set of definitions for the types of RDF streams present in the literature and current practice. Then, Section 3.2 organizes these definitions in the RDF Stream Taxonomy, and Section 3.3 links the proposed stream types to the state-of-the-art review performed in the previous section. The basis for all following definitions is the RDF 1.1 W3C Recommendation Cyganiak et al. (2014), which formally defines terms such as RDF graphs and RDF triples, and must be implemented by any RDF 1.1-compliant software. We use it with the aim of making our definitions readily applicable in the context of existing, RDF 1.1-compliant software.
The model-theoric semantics of RDF streams are not considered in this work. Instead, the proposed semi-formal definitions can be used as a universal basis on top of which the formal semantics can be defined, as needed. This is motivated by, firstly, the semantics of RDF streams being entirely irrelevant for some of the discussed application areas (e.g., streaming protocols, streaming I/O). Secondly, only the semantics of RDF graphs are well-defined – there are no agreed-upon formal semantics for RDF datasets Zimmermann (2014), and there are similarly many ways to approach this topic for RDF streams. Therefore, we leave this topic for future discussion. The definitions are formulated in a semi-formal manner similar to the RDF 1.1 Concepts and Abstract Syntax W3C Recommendation Cyganiak et al. (2014) or the RSP Data Model RDF Stream Processing Community Group , with the goal of making them readily applicable for RDF practitioners.
@@ 3.1 RDF Stream Definitions
{Definition}
An RDF stream is an ordered, potentially infinite sequence of RDF stream elements. The manner in which the sequence is ordered is not specified on purpose, as this depends on the use case. Often, the order will be defined by time, but a stream does not have to be time-dependent (in fact, many methods do not consider time at all – see Section 2). This definition also does not specify what exactly are the RDF stream elements, as this is left to the further definitions that are based on it. Thus, we specialize the definition to specify this:
{Definition}
@@
{Definition}
A flat RDF stream is an RDF stream whose elements are statements (either RDF triples or RDF quads). Grouped and flat RDF streams are still abstract, in the sense that their definitions are not concretized enough for practical implementation. However, they correspond to an important distinction between streams of groups of statements and streams of individual statements. With this, we can define the concrete types of RDF streams which can be used in practice:
{Definition}
An RDF graph stream is a grouped RDF stream whose elements are unnamed (default) RDF graphs. In other words, an element of an RDF graph stream is a set of triples. This is a commonly used formulation that can be further specialized:
{Definition}
An RDF subject graph stream is an RDF graph stream in which every element contains an IRI node (called the subject node) that uniquely identifies the graph in the stream. Every other node in the graph can be reached by traversing triples, starting from the subject node. RDF subject graph streams or similar formulations were proposed to handle timestamped RDF streams (TA-RDF Rodrıguez et al. (2009) and Stream Containers Schraudner and Harth (2022)). ERI Fernández et al. (2014) uses a similar formulation for subject-molecules, but its subject IRIs are unique only within one block. The proposed definition is slightly broader than the timestamped models, while still being useful – the advantage is that the element can be identified by an IRI, without using RDF datasets. The rest of the element can be discovered with what can be intuitively thought of as RDF graph connectivity.
@@
{Definition}
An RDF dataset stream is a grouped RDF stream whose elements are RDF datasets. It should be noted here that although an RDF graph stream may seem like a special case of an RDF dataset stream, it is not. According to RDF 1.1, an RDF dataset is not a generalization of an RDF graph, but rather a collection of RDF graphs. Thus, the definitions of RDF dataset streams and RDF graph streams are not directly related in the taxonomy. It should be noted, however, that an RDF graph stream can be trivially transformed into an RDF dataset stream, by simply assuming that each graph in the former is the default graph in the dataset.
@@
{Definition}
An RDF named graph stream is an RDF dataset stream in which every element has exactly one named RDF graph pair $\langle n,G\rangle$, where $G$ is an RDF graph, and $n$ is the graph name. Apart from graph $G$, the dataset may contain any number of triples in the default graph. This is a simplified version of the RSP Data Model RDF Stream Processing Community Group , without the temporal aspect. Although we could find only one reference to this stream type in the literature (RSP engines can consume such streams if an appropriate conversion is implemented RDF Stream Processing Community Group ), it is an intuitively important specialization of the RDF dataset stream. Here, every element corresponds to one named RDF graph with some optional information (metadata) about it in the default graph. As with RDF subject graph streams, this has the advantage of identifying a stream element by IRI, albeit with more flexibility as to the element’s contents.
@@
{Definition}
A timestamped RDF named graph stream is an RDF named graph stream in which every element is a timestamped named graph. The elements that share the same timestamp predicate $p$ are ordered by the partial order associated with $p$. Definitions 3.1 and 3.1 were derived directly from the draft RSP Data Model RDF Stream Processing Community Group . Finally, we define concrete flat RDF stream types:
{Definition}
@@
{Definition}
A flat RDF quad stream is a flat RDF stream whose elements are quads. An “RDF quad” is a term that is often used by practitioners, but not defined in the RDF 1.1 Recommendation. It is however defined in the RDF 1.2 Working Draft (16 April 2024) Hartig et al. (2024), and thus we use the RDF 1.2 draft definition of quads here.
@@
Grouped and flat RDF streams are notably interrelated, as any RDF graph stream can be flattened into an RDF triple stream, by simply listing and concatenating the statements in each of its elements – the same is true for RDF dataset streams and RDF quad streams. The reverse is also possible, the elements of a flat stream can be grouped into a grouped stream. Such operations are in fact common in streaming applications Bonte and Tommasini (2023); Fernández et al. (2014); Keskisärkkä and Blomqvist (2013); Sowiński et al. (2022).
Finally, for all of the above definitions, their generalized variants can be trivially derived (e.g., generalized RDF graph stream), by using generalized triples and generalized datasets as their basis. We skip these definitions for the sake of brevity.
@@ 3.2 RDF Stream Taxonomy
The proposed definitions form a taxonomical structure of task formulations (RDF Stream Taxonomy – RDF-STaX), summarized in Figure 1. This overview allows us to observe that each concrete RDF stream type uses one of the four element types: graph, dataset, triple, and quad, possibly with additional restrictions on them. In fact, these four types of elements seem to be the only conceivable ones that can be derived from the RDF 1.1 specification. For each of them, representing the stream element is a known and solved problem (RDF serializations), and thus they can be easily used in practical applications. The abstract stream types are not directly usable in practice by themselves, but rather serve as the basis of the taxonomy and a shared starting point for other definitions.
@@
3.3 Taxonomy Correspondence to the State of the Art
Finally, we link the proposed RDF stream types to the definitions used in state-of-the-art research and software – Table 1 presents such a summary. As noted in Section 2.6, a single RDF stream can be viewed from multiple perspectives, which is reflected in the table, with some items appearing more than once when such a situation occurs. In some cases the correspondence is not ideal and there may be insignificant differences between the definitions proposed here and those in the literature. The short descriptions presented in the table are expanded upon in the nanopublications dataset, as described in Section 5.3.
@@ 3.3 Taxonomy Correspondence to the State of the Art
RDF graph stream ActivityPub Tallon and Webber (2018), DBpedia-Live Morsey et al. (2012), ERI (blocks) Fernández et al. (2014), GoT Le-Phuoc et al. (2016), IPSM (logical level) Ganzha et al. (2017), Jelly (stream frames) Sowiński et al. (2022), Keskisärkkä & Blomqvist Keskisärkkä and Blomqvist (2013), Live open scientific knowledge graphs Le-Tuan et al. (2022), RDF EXI Käbisch et al. (2015), RDSZ Fernández et al. (2014), RSP Data Model (can be consumed) RDF Stream Processing Community Group , RSP4J YASPER (can be consumed) Tommasini et al. (2021), S-HDT Hasemann et al. (2012), VoCaLS Tommasini et al. (2018)
@@
ERI (subject-molecule stream, only within a block) Fernández et al. (2014), LDES Van Lancker et al. (2021), TA-RDF Rodrıguez et al. (2009), Stream Containers Schraudner and Harth (2022)
RDF dataset stream IPSM (output) Ganzha et al. (2017), Jelly (stream frames) Sowiński et al. (2022), RMLStreamer Haesendonck et al. (2019); Oo et al. (2022), RSP Data Model (can be consumed) RDF Stream Processing Community Group , Semantic Annotation enabler Szmeja (2023)
@@
RSP Data Model (can be consumed) RDF Stream Processing Community Group
@@
Barbieri & Della Valle Barbieri and Valle (2010), Linked Data Notifications for RDF streams Calbimonte (2017), RSP Data Model RDF Stream Processing Community Group , Tappolet & Bernstein Tappolet and Bernstein (2009), TripleWave Mauri et al. (2016), Wu et al. Wu et al. (2022)
@@
Apache Jena RIOT Apache Software Foundation (2023), ERI (high level) Fernández et al. (2014), Groppe et al. Groppe et al. (2007), Jelly (high level) Sowiński et al. (2022), Keskisärkkä & Blomqvist (mentioned) Keskisärkkä and Blomqvist (2013), RDF4J Rio Eclipse Foundation, Inc. (2023), RSP4J YASPER (can consume) Tommasini et al. (2021), SPARQL-Generate Lefrançois et al. (2016), VoCaLS Tommasini et al. (2018)
@@
Apache Jena RIOT Apache Software Foundation (2023), CARML CARML contributors (2023), Jelly (high level) Sowiński et al. (2022), RDF4J Rio Eclipse Foundation, Inc. (2023)
       The RSP Data Model states that these stream types can be consumed by RSP engines by adding timestamps to each stream element. This is an optional feature of the RSP Data Model.
The used representation differs from the definition in RDF-STaX by placing the metadata of the elements not in the default graph of the RDF dataset, but in a dedicated named graph.
@@ 4 RDF-STaX Ontology
To make the proposed taxonomy readily applicable, it was modeled as an OWL 2 DL ontology (Figure 2), available at https://w3id.org/stax/1.1.1/ontology (accessed on 4 June 2024) under the Creative Commons Attribution 4.0 license. In the ontology, RDF stream types are instances of either the stax:AbstractStreamType or the stax:ConcreteStreamType class, with taxonomical relations being realized using SKOS in the OWL DL version Jupp et al. (2008); Miles et al. (2005). The lightest ontology profile that could be used for RDF-STaX is the unofficial OWL 2 DL, due to RDF-STaX being based on SKOS. This balances usefulness in the contexts where ontological reasoning is used (by being compatible with a number of reasoners Singh et al. (2020)), with uses where ease of annotation takes precedence over semantics (e.g., in stream metadata or knowledge graphs with very little enforced semantics).
Each stream type instance in the RDF-STaX ontology is accompanied by a label, description, formal definition, and usage example. The instances are not shown in Figure 2, but the structure they form is identical to the one presented in the taxonomy overview diagram (Figure 1), realized using the skos:broader property.
@@
It should be stressed that RDF-STaX does not define a class for RDF streams, but rather classes for RDF stream types. For example, stax:namedGraphStream is an instance of the stax:ConcreteRdfStreamType class. The ontology also defines a class for RDF stream type usages (stax:RdfStreamTypeUsage), which can be thought of as subjective views of how an RDF stream is used in a specific context. This design philosophy is closely reflected in the expected usage patterns, described in the next subsection.
The ontology includes rich semantic relations between stream types. Properties stax:canBeFlattenedInto and stax:canBeGroupedInto relate grouped and flat stream types that can be easily converted into one another. For example, an RDF graph stream can be flattened into a flat RDF triple stream by concatenating its elements, while a flat RDF quad stream can be grouped into a RDF dataset stream by punctuating it. The property stax:canBeTriviallyExtendedInto is used for graph- and triple-based streams, to indicate that they can be trivially turned into dataset- and quad-based streams, respectively. This transformation consists of stating explicitly that the triples are in the default graph. Object property chains are defined in the ontology to allow reasoning on these relations, exploiting the taxonomical relations from SKOS Miles et al. (2005). These reasoning rules are presented below, in Manchester OWL Syntax Horridge et al. (2006):
{minted}
[breaklines]turtle skos:broaderTransitive o stax:hasElementType SubPropertyOf: stax:hasElementType
skos:broaderTransitive o stax:canBeFlattenedInto SubPropertyOf: stax:canBeFlattenedInto
skos:broaderTransitive o stax:canBeGroupedInto SubPropertyOf: stax:canBeGroupedInto
skos:broaderTransitive o stax:canBeTriviallyExtendedInto SubPropertyOf: stax:canBeTriviallyExtendedInto
The ontology in version 1.1.1 includes 5 classes, 8 object properties, 15 individuals, and a total of 192 axioms. The axiom count was obtained with the measure command in ROBOT 1.9.6 Jackson et al. (2019) and does not include the imported SKOS vocabulary. The ontology’s documentation (accessed on 4 June 2024) contains a detailed description of every class, property, and individual. Additional information facilitating ontology reuse is included, such as an explanation of the available ontology formats, download links, and versioning information.
4.1 Ontology Usage Patterns
The proposed usage patterns for the RDF-STaX ontology reflect the philosophy of having a subjective view of how an RDF stream is used. These patterns are described in detail in the ontology’s documentation (accessed on 4 June 2024), along with several illustrative examples and links to further resources.
The primary usage pattern in RDF-STaX is to create instances of the stax:RdfStreamTypeUsage class to annotate research works, software, or datasets, with each instance corresponding to a different view or use case of the stream. This pattern was selected due to the inherent subjectivity of determining the types of RDF streams, as discussed in Section 2.6. Below is an example of how this pattern may be used to annotate a DCAT Browning et al. (2023) Dataset:
{minted}